OpenStudy (anonymous):

Can any one help with this? I am just stumped on this problem! plot each point and form the triangle ABC. Verify that the triangle is a right triangle. Find its area. A =(4, −3); B =(4, 1); C = (2, 1)

OpenStudy (anonymous):

thats it roughly find lengths of the 3 sides and Check that AC^2 = AB^2 + BC^2 so proving ABC is rt angled

OpenStudy (anonymous):

No negative x values here, but not that important

OpenStudy (anonymous):

oh - right - i messed up there1

OpenStudy (anonymous):

The distance formula is \[d=\sqrt{(y2-y1)+(x2-x1)}\] You don't need to find the square root though because you're just going to be squaring them again when using the Pythagorean theorem.

OpenStudy (anonymous):

for example BC^2 = 0^2 + (4-2)^2 = 4

OpenStudy (anonymous):

Thanks guys :)

OpenStudy (anonymous):